# Collaborative Perception Datasets for Autonomous Driving: A Review

### v1

@@ II-D V2X Datasets
CP empowered by V2X technology facilitates data sharing and joint environmental understanding through communication among diverse agents, including vehicles, infrastructure, pedestrians, and other IoT devices. This paradigm significantly improves perception robustness and driving safety in autonomous systems.
From the perspective of agent structure and heterogeneity, recent V2X datasets reveal an increasing trend towards modeling complex interactions among heterogeneous agents with varying sensing capabilities. Early works such as V2X-SIM [41] support basic cooperative tasks, yet remain limited in agent type diversity and environmental complexity. The DAIR-V2X series [39, 55, 56] significantly expands the agent spectrum by introducing bi-agent collaboration between vehicles and RoadSide Units (RSUs). These datasets illustrate asymmetric collaboration in heterogeneous settings, where infrastructure and mobile vehicles possess distinct sensor setups and motion patterns.
To capture multi-agent temporal dynamics, V2XPnP [91] offers sequential data involving multiple cooperative vehicles, facilitating trajectory prediction in temporally correlated environments. DOLPHINS [88] further enhances agent diversity by simulating tri-vehicle collaboration in adverse conditions such as rain and fog, which supports evaluation of perception robustness under extreme scenarios. V2X-Real [54] emphasizes large-scale data collection across multiple vehicles and infrastructures, reflecting real-world V2X deployments with rich agent-level interactions. Adver-City [50] and DeepAccident [51] introduce environmental challenges such as snow, low visibility, and accident-prone zones, where CP across various agents becomes critical. Meanwhile, V2X-Radar [57] and V2X-R [58] expand the sensor modalities by incorporating 4D radar into the V2X sensing framework, highlighting the potential of cross-agent, cross-modal fusion in enhancing perception fidelity.
Despite these advancements, V2X datasets still face challenges related to scalability, diversity, and task integration. Many datasets lack coverage of complex urban intersections or multi-agent scenarios with mixed mobility patterns. Additionally, existing benchmarks often address specific tasks in isolation, limiting their support for evaluating end-to-end CP systems across heterogeneous agents.
@@ II-E I2I Datasets
As V2X CP continues to evolve, I2I collaboration has emerged as a promising paradigm to address the limitations of single-agent roadside systems. While V2I, V2I and V2X datasets focus primarily on perception enhancement via VIC, their perception range is often constrained by occlusion, especially when large vehicles or urban structures obstruct the view. I2I CP addresses this issue by enabling multiple roadside infrastructure nodes to share and fuse data, thereby reducing blind spots and improving the robustness of perception in complex environments such as intersections and road corridors.
Two pioneering datasets exemplify the advancement of I2I CP. The RCooper [60] is the first real-world, large-scale roadside CP dataset. It includes two typical scenarios—intersection and corridor—with infrastructure agents equipped with multi-type LiDARs and cameras. RCooper provides high-quality annotations, demonstrating the effectiveness of cross-infrastructure cooperation. In contrast, the InScope dataset [59] specifically targets occlusion-aware perception by deploying LiDARs at multiple viewpoints. Additionally, InScope proposes a novel quantitative metric to evaluate anti-occlusion capabilities, providing a rigorous framework for I2I perception evaluation. V2XPnP [91] further expands the paradigm by integrating large-scale sequential data across V2X modes, particularly I2I, offering valuable support for complex CP tasks. Compared to other paradigms, I2I collaboration is more effective in mitigating occlusion in dense traffic environments. However, Current I2I datasets face challenges in generalization, data synchronization, and scene diversity, highlighting the need for scalable deployment strategies and standardized evaluation protocols.
@@
Simulation and real-world datasets serve complementary roles in AD research, differing in cost, source, annotation accuracy, realism, and applicability. Simulation datasets, generated through virtual modeling, offer low cost, high scalability, and automatic annotations, enabling the design of diverse scenarios, including rare or hazardous conditions. However, the simulation-to-reality gap limits their generalization capabilities. Real-world datasets, collected from actual environments, provide greater complexity and realism, reflecting sensor noise and dynamic interactions, making them essential for final model validation. Despite their reliability, they are costly and labor-intensive to acquire and annotate, often lacking scenario diversity. As a result, simulation datasets are preferred for algorithm development and pre-training, while real-world datasets remain critical for evaluation and deployment [36].
@@ III-A1 Simulation Platforms
Simulators are essential for CP research in AD, offering cost-effective, safe, and annotation-efficient alternatives to real-world data collection. They enable accurate generation of object attributes and automated annotations. Open-source platforms such as CARLA [99], SUMO [100], and AirSim [101] are widely used for data synthesis. CARLA supports diverse driving scenarios and multimodal sensors, including RGB, depth, radar, and LiDAR, but its LiDAR simulation has limitations in point cloud intensity and echo dropout modeling. SUMO provides vehicle and pedestrian trajectory simulation and can be co-simulated with CARLA to improve scene fidelity. The OpenCDA [102] framework integrates both platforms, offering standardized scenario management and benchmarks for perception and planning tasks.
@@ III-A2 Simulation Datasets
Simulation datasets serve as vital testbeds for CP research, offering controllable environments, automated annotations, and support for diverse or rare scenarios. Representative datasets are built on CARLA, SUMO, or LiDARsim platforms. COMAP [79] and OPV2V [40] leverage CARLA-SUMO co-simulation for multi-agent perception across intersections and urban scenes, supporting detection, segmentation, and trajectory prediction. Extensions such as OPV2V+ [80] and OPV2V-H [83] improve agent diversity and simulate asynchronous sensing. DOLPHINS [88] and Adver-City [50] incorporate adverse weather to evaluate sensor robustness. V2X-Sim [41], Multi-V2X [89], and DeepAccident [51] expand support for high-resolution sensors, multi-agent scheduling, and accident prediction. These datasets provide critical foundations for benchmarking CP algorithms and exploring Sim2Real transfer, communication modeling, and generalization across scenarios.
@@ III-B Real-World Datasets
Real-world CP datasets provide authentic sensor data and traffic interactions essential for AD research and model validation. DAIR-V2X [39] is the first large-scale multimodal real-world dataset focusing on V2I cooperation, while V2X-Seq [42] extends it with long-sequence data and traffic signal information. V2V4Real [53] is the first large-scale real-world multimodal V2V dataset, offering valuable support for Sim2Real studies despite limited CAV diversity. LUCOOP [81], collected with three CAVs using heterogeneous sensors, supports urban driving studies, while MARS [82] expands to four CAVs for improved stability and generalization. HoloVIC [85] and OTVIC [86] focus on V2I setups, with OTVIC simulating real-time vehicle-RSU communication to reflect dynamic transmission challenges in infrastructure-assisted perception.
@@ IV Cooperative perception Scenario Datasets
CP datasets can be categorized according to different traffic environments to address the perception needs of AD systems in various scenarios. In this section, we classify them into three types: intersection, urban street, and highway scenarios. Fig. 5 shows the distribution of representative datasets across these categories. This chapter discusses the key datasets in each scenario and their characteristics, providing essential data support for CP research.
@@ IV-A Intersection Datasets
Ko-PER [43] is one of the earliest datasets employing fixed LiDAR and cameras to monitor intersection dynamics. IPS300+ [74] improves spatial resolution with high-density 3D annotations, while LUMPI [76] introduces multi-view data under varying weather and lighting. TUMTraf-I [77] further targets annotation precision in adverse conditions, such as night and rain. HoloVIC [85] designs diverse holographic intersections to explore sensor placement strategies for CP. Several datasets incorporate intersection scenarios within broader V2X contexts. DAIR-V2X [39] supports V2I perception in urban crossings; DOLPHINS [88] includes intersection and T-junctions with weather variations; DeepAccident [51] focuses on accident prediction at intersections. V2XPnP [91] offers 163 intersection scenarios, supporting trajectory prediction and object tracking. These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion. However, challenges such as dynamic traffic variability, sensor blind zones remain. These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion.
@@ IV-B Urban Street Datasets
Urban street CP datasets span V2V, V2I, and V2X paradigms, supporting tasks such as object detection, tracking, and trajectory prediction. CityFlow [44] offers multi-camera tracking data across arterials and residential areas. LUCOOP [81] collects time-synchronized multimodal data from three Connected Autonomous Vehicles (CAVs), including 3D point clouds, city models, and UWB ranging data. MARS [82] further extends to four CAVs under diverse weather, enhancing perception stability and agent collaboration. Several datasets strengthen roadside sensing capabilities. RCooper [60] provides the first large-scale real-world roadside dataset for multi-sensor infrastructure perception. RoScenes [78] delivers dense multi-view 3D data in complex urban traffic, supporting BEV perception. V2XPnP [91] captures diverse urban road scenes with multi-agent trajectories and detailed interactions. Together, these datasets address complex urban traffic environments, offering comprehensive data support for perception, fusion, and prediction in multi-agent V2X systems.
@@ IV-C Highway Datasets
With the growing deployment of ICVs on highways, CP datasets for highway scenarios have become increasingly important for traffic management and AD safety. Highway environments demand long-range detection, lane-level perception, and robust environmental sensing, especially under high-speed conditions where precise synchronization is critical.
@@ V Datasets by Sensor Modality
Sensor modality diversity plays a pivotal role in enhancing the robustness of CP systems, particularly in complex and dynamic environments [103][104]. As illustrated in Fig. 6, typical sensors employed in CP include cameras, LiDAR, and Radar, each corresponding to distinct perception modalities such as RGB images, dense 3D point clouds, and sparse Radar point clouds. The respective strengths and weaknesses of these sensors are summarized in Table IV, providing a comprehensive overview of their capabilities under varying operational conditions. This section categorizes existing CP datasets based on their acquisition modalities, including image-only, LiDAR-only, and multimodal datasets.
@@
Image-only datasets, while cost-effective and semantically rich, are limited by poor depth estimation. LiDAR-only datasets offer precise 3D localization but are constrained by high sensor cost and limited semantic information. Multimodal datasets, especially those combining camera and LiDAR, dominate recent research due to their complementary strengths. Moreover, Radar-integrated datasets have gained attention for their resilience to adverse weather, marking a growing trend in sensor fusion for reliable perception [57].
@@
• Performance degrades in rain, snow, and is affected by smoke, dust, and other weather conditions.
@@ V-A Image-only Datasets
Image-only datasets offer high-resolution visual data, supporting tasks such as object detection and semantic segmentation. However, the absence of depth information limits their performance in occlusion handling, distance estimation, and robustness under varying lighting. CityFlow [44] spans large-scale urban scenarios, supporting cross-camera tracking and vehicle ReID, with extensive spatial and camera coverage. WIBAM [105], captured from infrastructure-mounted cameras, focuses on 2D/3D object detection. RoScenes [78], the largest multi-view roadside dataset to date, targets vision-based BEV perception, addressing challenges in multi-camera fusion and dense traffic scenes. Despite their effectiveness in visual tasks, image-only datasets have limited utility in high-precision localization and CP.
@@ V-C Multi-modal Datasets
Multi-modal datasets, particularly those fusing LiDAR and cameras, are central to CP research. By combining LiDAR’s precise 3D spatial sensing with camera-derived semantic richness, these datasets improve environmental understanding and robustness [106][107]. Recently, the integration of radar has further enhanced perception under adverse weather and for long-range detection. This section summarizes key LiDAR-Camera and LiDAR-Radar-Camera datasets.
LiDAR-Camera datasets include both infrastructure-based and cooperative datasets. Fixed-infrastructure datasets such as Ko-PER [43], IPS300+ [74], Rope3D [75], A9-Dataset [73], and LUMPI [76] use roadside LiDAR and cameras to produce accurate 3D annotations. H-V2X [52], RCooper [60], and TUMTraf-I [77] extend perception to challenging conditions such as rain and night. UAV and traffic camera datasets like INTERACTION [45] and BAAI-VANJEE [108] support interaction modeling and driving behavior analysis. CP datasets such as OPV2V [40], IRV2V [98], V2V4Real [53], LUCOOP [81], and MARS [82] collect multi-agent LiDAR-Camera data in both real and simulated settings. OPV2V-N [97] enhances camera robustness, while V2V-QA [84] introduces LLM-based collaborative tasks. CoPerception-UAV [35] supplements V2V data with aerial views. In V2X contexts, DAIR-V2X-C [39], V2X-Seq [42], HoloVIC [85], and OTVIC [86] support multi-modal detection, tracking, and planning. Simulated datasets such as V2X-Sim 2.0 [41], V2XSet [34], DOLPHINS [88], DeepAccident [51], V2X-Real [54], and Multi-V2X [89] further address diverse tasks from accident prediction to V2X data augmentation. LiDAR-Radar-Camera datasets such as T & J [46], OPV2V+ [80], OPV2V-H [83], and V2X-Radar [57] enhance CP under adverse conditions by leveraging radar’s robustness to weather and long-range sensing. In summary, LiDAR-Camera datasets remain the cornerstone of CP research, while LiDAR-Radar-Camera fusion introduces new capabilities for perception in challenging environments.

### v2

@@ II-D V2X Datasets
V2X-based CP enables rich environmental understanding by linking vehicles, infrastructure, and other agents through communication. This paradigm enhances perception completeness and safety in dynamic traffic environments.
Recent V2X datasets increasingly emphasize agent diversity and heterogeneous collaboration. Early datasets like V2X-SIM [40] support basic CP tasks in structured settings. The DAIR-V2X series [35, 54] extends to V2I collaboration with asymmetric sensor setups. Multi-agent dynamics are explored in V2XPnP [92] and DOLPHINS [89], which incorporate sequential reasoning and adverse weather. Adver-City [49], DeepAccident [50], and V2X-Real [53] improve realism and environmental diversity, while V2XSet-Noise [58] injects simulated sensor and timing noise to evaluate robustness under deployment-like conditions, and V2X-Radar [56] incorporates 4D radar for cross-modal fusion.
Figure 7: Samples from representative V2X datasets. (a)(b)(c) V2VSet [33]; (d)(e)(f) DeepAccident [50]; (g)(h)(i) Adver-City [49].
As shown in Fig. 7, the three datasets reflect complementary directions in advancing V2X perception. V2VSet (a–c) offers clean synthetic environments for cooperative 3D detection and view fusion, but lacks dynamic risk modeling. DeepAccident (d–f) introduces accident-prone and motion-rich sequences to support forecasting tasks. Adver-City (g–i) emphasizes sensor degradation under fog, glare, and other adverse conditions. These samples illustrate a trade-off between controllability (V2VSet), temporal realism (DeepAccident), and visual robustness (Adver-City), jointly covering core challenges in V2X CP.
Despite progress, V2X datasets still face scalability and integration challenges. Many lack coverage of complex intersection layouts or hybrid traffic patterns. Benchmarks often isolate perception tasks, limiting support for end-to-end CP evaluation across heterogeneous agents and adverse conditions.
@@ II-E I2I Datasets
As V2X CP advances, I2I collaboration emerges as a key paradigm to address the limitations of single-agent roadside systems. While traditional V2I and V2X enhance perception via vehicle–infrastructure communication, they remain susceptible to occlusions from static objects or large vehicles. I2I mitigates this by linking multiple roadside perception units via low-latency channels, enabling real-time feature or decision-level fusion. This improves robustness in occlusion-prone scenes such as intersections and corridors, and supports scalable, decentralized cooperative frameworks.
Three recent datasets highlight advances in I2I perception. RCooper [60] is the first large-scale real-world roadside dataset, covering intersection and corridor scenes with heterogeneous infrastructure agents using diverse LiDARs and cameras. It provides 3D detection and tracking labels to validate cross-node fusion. InScope [59] emphasizes occlusion-aware perception using multi-view LiDARs and introduces a degradation metric for systematic evaluation. V2XPnP [92] complements these with long-sequence, multi-agent data—including I2I settings—to support temporal modeling and complex tasks.
Figure 8: Samples of I2I collaborative perception datasets. (a)–(d) RCooper [60]; (e) InScope [59].
Fig. 8 compares two I2I datasets through representative samples. RCooper (a)–(d) features multi-angle RGB images and LiDAR point clouds from distributed infrastructure nodes, enabling long-range coverage and blind spot reduction via inter-node fusion. In contrast, InScope (e) employs a dual-LiDAR setup for occlusion-aware perception in dense urban scenes. While RCooper emphasizes spatial coverage and corridor-scale extensiveness, InScope focuses on precise detection in visually constrained environments, supported by its anti-occlusion evaluation metric. These datasets reflect complementary strategies in I2I perception: broad area fusion versus localized occlusion resolution.
Compared to other paradigms, I2I collaboration proves particularly effective in mitigating occlusions within dense traffic environments. However, current I2I datasets still face limitations in generalization, data synchronization, and scene diversity, underscoring the need for scalable deployment strategies and standardized evaluation protocols. Beyond I2I, emerging collaborative paradigms such as Vehicle-to-Network (V2N), cloud-based collaboration, and Vehicle-to-Pedestrian (V2P) are drawing increasing attention. V2N and cloud communication facilitate global map sharing and large-scale model updates, while V2P extends perception capabilities to vulnerable road users, broadening the scope and societal impact of CP systems.
@@
While Section II organized CP datasets according to communication paradigms such as V2V, V2I, and V2X, this section shifts perspective to the origin and fidelity of the data—specifically, the distinction between simulated and real-world datasets. This transition enables a deeper understanding of how dataset generation methods impact the development, evaluation, and deployment of CP systems.
Simulation and real-world datasets play complementary roles in collaborative perception. Simulated datasets offer scalable, controllable environments with dense annotations, enabling rapid iteration and benchmarking. In contrast, real-world datasets provide authentic sensor artifacts, diverse environmental conditions, and deployment realism essential for evaluating generalization and robustness. However, the domain gap between synthetic and real data introduces significant challenges for transferability and real-world performance [62]. In this section, we first review mainstream simulation platforms and representative simulated datasets, then analyze widely used real-world datasets. We conclude by discussing the Sim2Real gap, a key challenge in CP research, where models trained in simulation often degrade when applied to real-world environments. Table III summarizes the core differences between simulated and real-world datasets, highlighting their respective advantages and limitations.
@@ III-A1 Simulation Platforms
Simulation platforms play a pivotal role in CP research for AD, providing scalable, controllable, and annotation-rich alternatives to costly real-world data collection. By enabling the generation of synthetic sensor data with fine-grained ground truth, these simulators accelerate the development and validation of perception algorithms under diverse and repeatable conditions.
Among open-source simulation platforms, CARLA [102] is one of the most widely adopted tools for AD research. It supports various multimodal sensors such as RGB cameras, LiDAR, radar, and depth maps, and provides configurable urban and rural scenarios. However, its LiDAR simulation has limitations in modeling realistic point cloud intensity and echo effects. SUMO [103] focuses on microscopic traffic flow and is often used to simulate vehicle and pedestrian trajectories. When integrated with CARLA, it enhances multi-agent interaction realism. To facilitate CP and V2X studies, OpenCDA [104] integrates both platforms, offering standardized scenarios, communication protocols, and benchmarks for end-to-end testing. AirSim [105] is another widely used simulator for vision-centric tasks, providing photorealistic environments, though it offers limited support for LiDAR and complex traffic scenes.
@@ III-A2 Simulation Datasets
Simulation datasets play a fundamental role in CP research by enabling controlled data generation, automatic annotation, and repeatable evaluation under diverse or rare conditions. Most simulated datasets are built upon open-source platforms such as CARLA, SUMO, and LiDARsim, which support flexible scenario modeling and multi-sensor simulation.
Among them, COMAP [80] and OPV2V [39] employ CARLA-SUMO co-simulation to create interactive, multi-agent urban driving environments. These datasets support key CP tasks, including 3D object detection, semantic segmentation, and trajectory prediction. The OPV2V series has been extended in OPV2V+ [81] and OPV2V-H [84], introducing increased agent diversity and modeling asynchronous perception to reflect real-world temporal misalignment.
Several recent datasets focus on robustness under challenging environmental conditions. DOLPHINS [89] and Adver-City [49] simulate adverse weather scenarios such as rain, fog, and snow, which are critical for evaluating sensor degradation. In parallel, V2X-Sim [40], Multi-V2X [90], and DeepAccident [50] introduce large-scale cooperative settings with support for high-resolution sensor streams, complex agent coordination, and accident scenario modeling. Together, these datasets provide robust testbeds for benchmarking CP models, studying Sim2Real transfer, and analyzing communication-aware perception strategies.
@@ III-B Real-World Datasets
Real-world datasets are essential for validating CP systems under authentic sensor noise, traffic dynamics, and environmental complexity. DAIR-V2X [35] is the first large-scale multimodal real-world dataset focusing on V2I cooperation, while V2X-Seq [41] extends it with long-sequence data and traffic signal information. V2V4Real [52] is the first large-scale real-world multimodal V2V dataset, offering valuable support for Sim2Real studies despite limited CAV diversity. LUCOOP [82], collected with three CAVs using heterogeneous sensors, supports urban driving studies, while MARS [83] expands to four CAVs for improved stability and generalization. HoloVIC [86] and OTVIC [87] focus on V2I setups, with OTVIC simulating real-time vehicle-RSU communication to reflect dynamic transmission challenges in infrastructure-assisted perception. These datasets collectively bridge the gap between simulated environments and field deployments, fostering more transferable and robust CP systems.
III-C Sim2Real Gap
The Sim2Real gap refers to the performance degradation observed when models trained on synthetic data are deployed in real-world environments. This challenge is especially pronounced in CP, where heterogeneous sensors, complex traffic dynamics, and multi-agent coordination amplify domain discrepancies. Although simulation platforms offer controllable and annotation-rich environments, their sensor data often lacks realistic noise, occlusions, and temporal misalignment. Real-world deployments, in contrast, introduce GPS drift, communication latency, and inconsistent sensor configurations, resulting in both geometric and semantic mismatches that undermine model generalization.
To study this problem, several benchmark datasets have been developed. Simulated datasets such as OPV2V [39] and V2XSet [33] enable multi-agent experiments but fall short in capturing real-world uncertainties. Real-world datasets like V2V4Real [52] and DAIR-V2X [35] offer more realistic settings with diverse sensor setups. Empirical studies have quantified the Sim2Real gap across these datasets: S2R-ViT [106] reports up to 40% AP drop when transferring from OPV2V to V2V4Real under deployment noise; CoPEFT [107] improves AP@0.7 from 21.7% to 41.8% on DAIR-V2X via parameter-efficient tuning. DUSA [108] and CUDA-X [109] achieve 5–20% AP gains through domain-alignment modules. V2X-ReaLO [110] further reveals a 13.2 mAP drop during online deployment, highlighting the impact of latency and pose noise.
Collectively, these findings show that the Sim2Real gap in CP stems from both data distribution shifts and system-level constraints. Addressing it requires more realistic simulation environments, domain adaptation techniques, and hybrid training protocols that bridge synthetic and real domains. Future research should also prioritize time-aware, uncertainty-aware modeling and develop standardized benchmarks for cross-domain evaluation to ensure robust real-world deployment.
@@ IV Cooperative perception Scenario Datasets
Building on the discussion of data sources in Section III, we now focus on the deployment contexts of CP datasets. Scenario-specific datasets are essential for evaluating system performance under diverse traffic conditions and supporting targeted algorithm development. This section categorizes CP datasets into three representative scenarios: intersections, urban streets, and highways. These settings differ in traffic complexity, agent density, and occlusion patterns, shaping dataset design and system requirements. Fig. 9 illustrates the distribution of representative datasets. For each category, we review key datasets, emphasizing their design motivations, sensor configurations, and task relevance.
@@ IV-A Intersection Datasets
Ko-PER [42] is one of the earliest datasets using fixed LiDAR and cameras to observe intersection dynamics. IPS300+ [75] enhances resolution with dense 3D annotations, while LUMPI [77] introduces multi-view data under diverse weather and lighting. TUMTraf-I [78] focuses on annotation precision in night and rain conditions. HoloVIC [86] designs holographic intersections to study sensor deployment. Several V2X datasets also incorporate intersection scenarios: DAIR-V2X [35] supports V2I in urban crossings; DOLPHINS [89] adds weather-varied intersections; DeepAccident [50] centers on accident prediction; V2XPnP [92] includes 163 intersections for trajectory prediction and tracking. These datasets emphasize high-precision perception in occlusion-prone intersections via multi-view perception and cross-agent fusion, though challenges like traffic variability and sensor blind spots remain.
@@ IV-B Urban Street Datasets
Urban street CP datasets span V2V, V2I, and V2X paradigms, enabling tasks such as object detection, tracking, and trajectory prediction. CityFlow [43] provides multi-camera tracking across arterial and residential roads. LUCOOP [82] uses synchronized multimodal data from three Connected and Autonomous Vehicles (CAVs), including 3D point clouds, city models, and UWB ranging. MARS [83] extends to four CAVs under diverse weather, enhancing collaborative robustness. Roadside-focused datasets also emerge: RCooper [60] is the first large-scale real-world roadside CP dataset; RoScenes [79] offers dense multi-view 3D urban scenes for BEV perception; V2XPnP [92] captures multi-agent urban trajectories and interactions. Collectively, these datasets provide comprehensive urban traffic data for multi-agent perception, fusion, and prediction.
@@ IV-C Highway Datasets
With the growing deployment of ICVs on highways, CP datasets for highway scenarios have become increasingly important for traffic management and AD safety. Highway environments demand long-range detection, lane-level perception, and robust environmental perception, especially under high-speed conditions where precise synchronization is critical.
@@ V Datasets by Sensor Modality
Sensor modality diversity is critical for enhancing the robustness of CP systems in complex, dynamic environments [111][112]. As shown in Fig. 10, CP typically employs cameras, LiDAR, and Radar, corresponding to RGB images, dense 3D point clouds, and sparse Radar reflections. Image-only datasets are cost-effective and semantically rich but suffer from limited depth perception and sensitivity to lighting. LiDAR-only datasets provide accurate 3D localization and perform well in low-light conditions but lack semantic information and are costly to deploy. Multimodal datasets, particularly those integrating cameras and LiDAR, are prevalent due to their complementary strengths. Radar-integrated datasets have gained traction for their robustness in adverse weather [56]. Table IV summarizes the trade-offs among these modalities in terms of resolution, cost, weather resilience, and perception reliability. This section categorizes CP datasets based on sensor modality into image-only, LiDAR-only, and multimodal types.
@@
• Performance degrades in rain, snow, and affected by smoke, dust, and weather conditions.
@@ V-A Image-only Datasets
Image-only datasets offer high-resolution visual data, supporting tasks such as object detection and semantic segmentation. However, the absence of depth information limits their performance in occlusion handling, distance estimation, and robustness under varying lighting. CityFlow [43] spans large-scale urban scenarios, supporting cross-camera tracking and vehicle ReID, with extensive spatial and camera coverage. WIBAM [113], captured from infrastructure-mounted cameras, focuses on 2D/3D object detection. RoScenes [79], the largest multi-view roadside dataset to date, objects vision-based BEV perception, addressing challenges in multi-camera fusion and dense traffic scenes. Despite their effectiveness in visual tasks, image-only datasets have limited utility in high-precision localization and CP.
@@ V-C Multi-modal Datasets
Multi-modal datasets, particularly those fusing LiDAR and cameras, are central to CP research. By combining LiDAR’s precise 3D spatial perception with camera-derived semantic richness, these datasets improve environmental understanding and robustness [114][115]. Recently, the integration of radar has further enhanced perception under adverse weather and for long-range detection. This section summarizes key LiDAR-Camera and LiDAR-Radar-Camera datasets.
LiDAR-Camera datasets are foundational to CP research, covering both infrastructure-based and collaborative scenarios. Ko-PER [42], IPS300+ [75], Rope3D [76], and A9-Dataset [74] use roadside sensors for accurate 3D annotation, while H-V2X [51], RCooper [60], and TUMTraf-I [78] address challenging conditions such as rain and nighttime. INTERACTION [44] supports behavioral analysis from UAV perspectives.
Collaborative datasets like OPV2V [39], IRV2V [100], and V2V4Real [52] enable multi-agent perception. OPV2V-N [99] improves image robustness, and V2V-QA [85] introduces LLM-based reasoning. In V2X contexts, DAIR-V2X-C [35], V2X-Seq [41], and HoloVIC [86] support detection and planning. Simulated datasets such as V2X-Sim 2.0 [40], V2XSet [33], DOLPHINS [89], DeepAccident [50], V2X-Real [53], and Multi-V2X [90] address a wide range of cooperative tasks.
LiDAR-Radar-Camera datasets, including OPV2V+ [81], OPV2V-H [84], and V2X-Radar [56], enhance robustness under adverse conditions. Despite progress, challenges remain in aligning heterogeneous data and ensuring scalability across modalities—future work should emphasize cross-domain generalization and efficient multimodal fusion strategies.